# Multi-Agents are Social Groups: Investigating Social Influence of Multiple Agents in Human-Agent Interactions

### v1

@@ 4.3.2. Education Level
When analyzing the linear regressions between education level and DVs (normative influence, opinion change), the results suggest a strong relationship between education level and normative influence in 1-agent condition (Topic 2: $\beta$ = -0.2625, $p$ 0.05, $R^{2}$ = 0.132; Different Stance: $\beta$ = -0.3695, $p$ 0.01, $R^{2}$ = 0.236). While in 3-agent and 5-agent, the trend were not observed. This suggests that participants with a lower education level tended to perceived more normative influence compared to participants with a higher education level only with a single agent; however, with multiple agents, all participants were similarly affected. As for opinion change, no significant differences were found across the three conditions.
@@ 5.1. Multi-Agent Systems as Social Groups
Overall, our study demonstrates that virtual agents can function as a cohesive social group, leading users to experience social pressure. This mirroring of human group dynamics echoes the well-known CASA (Computers as Social Actors) paradigm (Nass et al., 1994), while extending it beyond its traditional focus on single agent systems. We propose that beyond being perceived as individual actors, systems of multiple agents can act as social groups, exhibiting many of the same social influences and pressures as groups of humans.
These social influence effects are more pronounced in some cases - for instance, when multiple agents were arguing against the user’s stance rather than for it. Multi-agent systems hence show potential in effecting attitude and opinion change more effectively than single-agent systems. However, there are some considerations that need to be taken into account when designing such systems.
@@
An interesting phenomenon emerges when comparing the results of perceived social influence and opinion change. In both cases, these effects were observed more with the multi-agent groups compared to the single-agent group. We generally saw participants in the 5-agent group reporting more pressure to agree with the agents than those in the 1-agent group; yet, with opinion change, it was those in the 3-agent group who had a significantly greater shift in opinion.
We suggest an explanation for this relation: the strongly one-sided dynamic of the discussion in the 5-agent group caused the significantly increased sense of pressure, in turn eliciting a negative reaction in some participants making them resistant to the agents’ arguments - a phenomenon known as psychological reactance (Steindl et al., 2015). Reactance theory describes threats to one’s freedom to act or think in a certain way - such as social influence targeting a particular individual (Clee and Wicklund, 1980) - and how they affect subsequent behavior. The increased sense of threat in the 5-agent group can be seen in the greater measured social pressure, and is further borne out by the responses to the question on perceived pressure, where stronger phrases such as ”extremely pushy” (P77), ”ganged up on” (P77, P79), and ”felt cornered” (P86) were found. Reactance theory suggests that threats ”serve as a motivator to restore one’s freedom”, with a core reaction to being threatened being to directly ”exhibit the restricted behavior” (Steindl et al., 2015) - in this case, the stance or opinion that each participant holds, yet is being persuaded against by the agents. We see evidence of this in the number of participants who shifted their opinions away from that of the agents: 12 participants (30%) in the 5-agent group, as compared to 7 participants (21%) in the 3-agent group. Among the 12 participants in the 5-agent group, sentiments include feeling like the agents ”all had an agenda” (P74), were in ”a high school clique I wasn’t apart [sic] of” (P77), ”ganged up on me” (P79), and ”cornered” the participant (P86). This then highlights an interesting dynamic in human-agent interaction: not unlike what might be seen in human groups, overt amounts of pressure from other agents to change one’s opinion can create a reactionary effect of rejecting the change and holding on more strongly to the initial opinion instead.
@@ 5.2. AI Agents Creating Social Norms
An interesting phenomenon arises as a natural result of our experimental setup. As the agents in the 3- and 5-agent group expressed their views, a social norm begins to form among the agents: their repeated endorsement of a particular point of view creates (the perception of) a shared understanding of what the correct opinion is (Chung and Rimal, 2016). There are a great many types of social norms which we do not investigate further; however, the social norm formed here is unique in one particular aspect. Most systems that utilize social norms as a means of persuasion convey computer-mediated social norms: behaviors or attitudes of a group of people that are communicated to the user via social norm messaging (Bonan et al., 2020; Snow et al., 2020, 2021; Chiu et al., 2020) (e.g. ”50% of people recycle their waste”). In contrast, our setup introduces the unique idea of computer-created social norms. During the conversation, the agents’ opinions were expressed as their own (e.g. ”I think self-driving cars are good”), rather than being relayed from humans. [^4] While some existing systems may use agents that express opinions as their own, these single agent setups lack the shared expectations, accepted behaviors, and collective agreement that are characteristic of social norms (Cialdini, 1998). With multiple agents echoing one another, this created a sense of collective agreement characteristic of a social norm. Even in the short span of their interaction with the agents, there were hints of participants feeling the collective agreement of the agents, with P86 calling them ”like-minded” and P88 feeling they ”seemed like a bandwagon”. This potential ability to craft a social norm and establish it using a multi-agent setup opens up many opportunities for designers of persuasive systems.
@@ 5.3.1. Persuasive Intervention Systems
Many current health interventions, both physical and digital (e.g. (Balloccu et al., 2021; Oyebode et al., 2021)), rely on persuasive means to nudge people toward healthier choices regarding harmful behaviors like smoking or excessive drinking. These interventions may use social norm messaging (e.g. ”Most people do not smoke” or ”On average, people believe smoking is harmful”) or personalized messages or coaching (e.g. through interactive interfaces), aiming to shift perceptions and encourage behavior change. However, while these approaches often rely on the concept of social persuasion, they neglect the core factor of group size in creating social influence. Our findings suggest that designers in fields such as health and and wellness technology can leverage multi-agent interfaces to create social norms and influence, enhancing the effectiveness of their systems. As in our study, without needing to create additional content, simply increasing the number of perceived virtual presences in a system may increase the effectiveness of its messaging. For instance, an interactive chatbot could adopt multiple personas, while a health or diet tracker app could introduce different agents that appear at different points throughout the app. By creating a cohesive group of agents that reinforce a shared stance, these systems can exert subtle social pressure that more effectively conveys a message or nudges a user towards complying with certain actions to achieve their goal.
@@
In our study, we designed our 3- and 5-group agents to exhibit a sense of cohesion and unity. The agents occasionally reaffirmed one anothers’ viewpoints (e.g. ”Those are valid points, Bella”, ”I couldn’t agree with Cody more”) and engaged with other agents’ arguments (e.g. ”That’s a thought-provoking point, Cody”). That these 3- and 5-agent groups were successful in creating social influence aligns with well-established literature, that normative social influence is ”greater among individuals forming a group than among an aggregation of individuals who do not compose a group” (Deutsch and Gerard, 1955). We therefore suggest that agents in a multi-agent system should exhibit unity and cohesion to be perceived as a group, in order to create social influence.
We note, however, that we do not suggest that a group of agents that do not verbally reinforce one anothers’ views cannot also exert social influence. Group cohesion or similarity can manifest in many ways, such as appearance, language style, identity, and more. We leave this for future work.
@@ 5.3.3. Optimal Number of Agents for Opinion Shifts
In designing multi-agent systems, our findings show that simply increasing the number of agents in a system does not necessarily help opinion change - indeed, increasing from three to five agents had the reverse effect of inducing psychological reactance and reducing this change. While the correlation between group size and social pressure is well-known (Asch, 1955), there is scarce literature on the concrete threshold of group numbers needed to trigger psychological reactance (even in traditional human psychology). More research is needed to investigate how many, and under what conditions, a given number of agents will or will not result in resistance rather than conformance to a suggested opinion.
However, regardless of group size, our study suggests the importance of avoiding designs where users feel ”ganged up” upon by the agents. Even though in our study, only the 5-agent group participants clearly reported these feelings, there were also hints of such feelings in the 3-agent group: P51 felt like an outsider ”for a split second”, and P57 ”kind of felt like the odd one out”. Designers should therefore take care to avoid inducing such negative feelings; possible ways to do so include having the agents express empathy and practice active listening, or having a small portion of the bots side with instead of against the user (a phenomenon similar to minority influence (Gardikiotis, 2011)). These are promising directions for future work.
@@ 5.4. Limitations & Future Work
First, given the exploratory and foundational nature of our study, we chose contemporary social issues that we considered to be fairly neutral, yet relevant enough in public discourse that participants would be capable of discussing them with the agents. With our study demonstrating the presence of social influence effects in these issues, future studies could explore potentially more contentious areas like health, ethics, or politics. Given that people tend to hold stronger and possibly more polarized opinions on these issues, future work could explore if multi-agent social influence effects still persist in such cases, or whether new effects might emerge.
Secondly, our study was conducted as a one-time experiment lasting less than an hour, which may not fully capture long-term attitude changes. People’s opinions may shift more or less over time, depending on various factors. Future research could incorporate longer interventions and follow-up measurements to explore how multi-agent influence evolves over time.
Thirdly, since social influence can be shaped by cultural norms, the cultural background of participants is another important consideration. Our study included participants from predominantly Western, individualistic cultures. In more collectivist societies, where social norms may have a stronger influence, the results could differ. Future work could examine the impact of multi-agent systems on social influence across diverse cultural contexts.
@@ 6. Conclusion
As multi-agent systems become increasingly integrated into daily life, understanding how these agents interact with humans and influence our perceptions and opinions remains an underexplored yet significant issue. To address this gap, we investigated how interacting with varying numbers of agents affects individuals’ perceptions and opinions through a human-agent social discussion experiment. By analyzing quantitative and qualitative data, we found that multi-agent groups led to more substantial shifts in participants’ opinions towards those of the agents (RQ1). Additionally, participants reported a stronger sense of normative influence when interacting with multiple agents, especially when the agents disagreed with them. We identified potential factors contributing to this social influence, including the felt pressure to conform motivated by a desire to affiliate with the group of agents (RQ2). Furthermore, demographic analysis indicated that younger participants were more likely to be influenced by multi-agents, potentially due to their higher susceptibility to social pressure and norms compared to older individuals (RQ3). Overall, these findings suggest that agents can function as a social group, exerting social influence on human participants similar to that of human groups. This insight extends the existing CASA framework by demonstrating how interactions with multiple agents can create a sense of group identity. Our study also underscores the potential of multi-agent systems for behavior intervention applications, offering design recommendations while cautioning against the risks and potential harms associated with their misuse.
@@ A.1.1. Quantitative
The analysis methods used here are consistent with those introduced in Section 3.5.
Item Scenario 1 Agent 3 Agents 5 Agents
Mean STD Mean STD Mean STD F p
Understanding Topic1 5.35 1.84 5.45 1.06 5.30 1.37 1.28 0.53
@@ A.1.2. Qualitative
A.2. Survey Items
A.2.1. RQ1 - Opinion Change

### v2

@@ 4.3.2. Education Level
When analyzing the linear regressions between education level and DVs (normative influence, opinion change), the results suggest a strong relationship between education level and normative influence in 1-agent condition (Topic 2: $\beta$ = -0.2625, $p$ ¡ 0.05, $R^{2}$ = 0.132; Different Stance: $\beta$ = -0.3695, $p$ ¡ 0.01, $R^{2}$ = 0.236). While in the 3-agent and 5-agent, the trends were not observed. This suggests that participants with lower education level tended to perceive more normative influence compared to participants with higher education level in the 1-agent condition. As for opinion change, no significance was found across the three conditions.
@@ 5.1. Multi-Agent Systems as Social Groups
Overall, our study demonstrates that users were more likely to shift their opinions when agents expressed opposing views, and became more polarized when agents reinforced their initial stance. These patterns highlight that virtual agents can function as cohesive social groups, capable of exerting social pressure and evoking a desire for affiliation—key hallmarks of human group dynamics. This extends the classic CASA (Computers as Social Actors) paradigm (Nass et al., 1994), which primarily focused on single-agent interactions, by showing that users can also attribute group-like characteristics and influence to systems of multiple agents.
Importantly, while these behavioral outcomes resemble those observed in human-human interaction, our findings reveal a critical distinction in the underlying mechanisms of social influence. Specifically, multi-agent systems appear to elicit normative influence, where users conform to align with group norms or avoid social discord, but do not trigger informational influence in the same way human groups do. In human contexts, people often interpret consensus among others as a cue for informational validity or correctness (see Section 2.2). However, both our quantitative and qualitative results indicate that users did not perceive consensus among AI agents as making an opinion more accurate or trustworthy (see Section 4.2.1).
Based on prior work on public perceptions of AI systems (Luger and Sellen, 2016; Binns et al., 2018), this difference may stem from how users conceptualize the nature of AI agents. While human group members are typically seen as independent thinkers with diverse perspectives and life experiences, AI agents are often viewed as operating from the same underlying system or dataset (Luger and Sellen, 2016; Binns et al., 2018). As a result, when several agents present the same view, users may interpret it as repetition rather than meaningful reinforcement, increasing social pressure, but not enhancing perceived credibility or informativeness.
This distinction is important: it suggests that while AI agents can mimic group consensus behaviorally, they may not yet replicate the communicative richness and diversity that make human group influence persuasive. Future multi-agent systems could explore ways to design agents with varied backgrounds, perspectives, and communication styles to foster not just social pressure, but also perceived informational value and trust.
5.2. More Agents Are Not Always More Persuasive
@@
An interesting phenomenon emerges when comparing the results of perceived social influence and opinion change. As expected, both of these effects were observed more with the multi-agent groups compared to the single-agent group. In particular, participants in the 5-agent group reported more pressure to agree with the agents than those in the 1-agent group. However, with opinion change, it was those in the 3-agent group who had a significantly greater shift in opinion, while those in the 5-agent group did not.
We suggest that the contrasting patterns of opinion change and polarization can be understood through the lens of psychological reactance (McDonald and Crandall, 2015), which refers to individuals’ tendency to resist persuasion when they perceive it as a threat to their autonomy. This mechanism is particularly relevant in the different-stance condition, where agents challenge the participant’s original opinion. In this context, unanimous disagreement from five agents may have created an overtly strong sense of pressure (see Section 4.2.2). Psychological reactance theory (McDonald and Crandall, 2015) suggests this could have been perceived as a threat to their freedom of choice, triggering reactive resistance, in which participants rejected the agents’ stance, reinforcing their own opinion or even shifting away from the agent’s viewpoints in a bid to preserve their sense of autonomy and self-agency. By contrast, the 3-agent group may have conveyed enough social consensus to be persuasive without crossing the threshold that provokes reactance.
In the same-stance condition, however, agents supported the participant’s existing beliefs. In this case, without perceived conflict or coercion, psychological reactance is less likely to occur (Rosenberg and Siegel, 2018). As a result, a greater number of agreeing agents—such as in the 5-agent group—can amplify the perceived consensus and reinforce prior beliefs, leading to greater opinion polarization. This distinction helps explain our findings: opinion change was most prominent in the 3-agent condition under disagreement, while opinion polarization was most pronounced in the 5-agent condition under agreement.
Together, these results highlight how the persuasive effect of agent group size is shaped not only by the number of agents but by their alignment with the user’s initial stance, underscoring that more agents do not always yield stronger persuasive outcomes.
Future AI system designers could consider incorporating strategies known to reduce psychological reactance (Rosenberg and Siegel, 2018). Two commonly used approaches, preemptive messaging (e.g., ”You might not agree, and that’s okay—we just want to share some viewpoints.”) and post-message restoration (e.g., ”We respect your views—thank you for considering ours.”) (Richards et al., 2022), can be adapted to multi-agent settings to maintain persuasive effectiveness while preserving users’ sense of autonomy. By embedding these strategies into multi-agent system design, developers can reduce perceived coercion and enhance the system’s ability to persuade without triggering resistance, particularly in scenarios where disagreement is expected.
@@ 5.3. AI Agents Creating Social Norms
An interesting phenomenon arises as a natural result of our experimental setup. As the agents in the 3- and 5-agent group expressed their views, a social norm begins to form among the agents: their repeated endorsement of a particular point of view creates (the perception of) a shared understanding of what the correct opinion is (Chung and Rimal, 2016). There are many other types of social norms that we do not investigate further; however, the social norm formed here is unique in one particular aspect. Most systems that utilize social norms as a means of persuasion convey computer-mediated social norms: behaviors or attitudes of a group of people that are communicated to the user via social norm messaging (e.g. ”50% of people recycle their waste”) (Bonan et al., 2020; Snow et al., 2020, 2021; Chiu et al., 2020). In contrast, our setup introduces the unique idea of computer-created social norms. During the conversation, the agents’ opinions were expressed as their own (e.g., ”I think self-driving cars are good”), rather than being relayed from humans. [^4] While some existing systems may use agents that express opinions as their own, these single-agent setups lack the shared expectations, accepted behaviors, and collective agreement that are characteristic of social norms (Cialdini, 1998). With multiple agents echoing one another, this created a sense of collective agreement characteristic of a social norm. Even in the short span of their interaction with the agents, there were hints of participants feeling the collective agreement of the agents, with P86 calling them ”like-minded” and P88 feeling they ”seemed like a bandwagon”. This potential ability to craft a social norm and establish it using a multi-agent setup opens up many opportunities for designers of persuasive systems.
We note that we only studied agent groups in which all agents held the same opinion. This was a deliberate choice to investigate what we considered the most fundamental case of social norms, in which multiple individuals (or agents, in our case) with the same opinion create social pressure on others to change their opinion. This corresponds to the possibly realistic scenario of system designers attempting to engineer opinion change by replicating the multi-human social influence effect with a group of agents. Beyond this fundamental scenario, however, other group compositions such as one agent with a dissenting opinion or agents agreeing to different degrees with a given opinion have also been studied in human-human interactions; this presents a promising avenue for future work.
@@ 5.4.1. Persuasive Intervention Systems
Many current health interventions, both physical and digital (e.g. (Balloccu et al., 2021; Oyebode et al., 2021)), rely on persuasive means to nudge people toward healthier choices regarding harmful behaviors like smoking or excessive drinking. These interventions may use social norm messaging (e.g. ”Most people do not smoke” or ”On average, people believe smoking is harmful”) or personalized messages or coaching (e.g. through interactive interfaces), aiming to shift perceptions and encourage behavior change. However, while these approaches often rely on the concept of social persuasion, they neglect the core factor of group size in creating social influence. Our findings suggest that designers in fields such as health and and wellness technology can leverage multi-agent interfaces to create social norms and influence, enhancing the effectiveness of their systems (Lee et al., 2025). As in our study, without needing to create additional content, simply increasing the number of perceived virtual presences in a system may increase the effectiveness of its messaging.
Moreover, our results show that younger users and those with lower education levels are more susceptible to influence from multiple AI agents. This suggests that persuasive intervention systems targeting younger populations—such as those focusing on education (DeLay et al., 2016; Cox, 2023), prosocial behavior (Zhu et al., 2025), or adolescent smoking prevention (Kobus, 2003)—could particularly benefit from multi-agent designs. For instance, an interactive chatbot could adopt multiple personas, while a health or diet tracker app could introduce different agents that appear at different points throughout the app. By creating a cohesive group of agents that reinforce a shared stance, these systems can exert subtle social pressure that more effectively conveys a message or nudges a user towards complying with certain actions to achieve their goal.
@@
In our study, we designed our 3- and 5-group agents to exhibit a sense of cohesion and unity. The agents occasionally reaffirmed one anothers’ viewpoints (e.g., ”Those are valid points, Bella”, ”I couldn’t agree with Cody more”) and engaged with other agents’ arguments (e.g., ”That’s a thought-provoking point, Cody”). That these 3- and 5-agent groups were successful in creating social influence aligns with well-established literature, that normative social influence is ”greater among individuals forming a group than among an aggregation of individuals who do not compose a group” (Deutsch and Gerard, 1955). We therefore suggest that agents in a multi-agent system should exhibit unity and cohesion to be perceived as a group, in order to create social influence.
We note, however, that we do not suggest that a group of agents that do not verbally reinforce one another’s views cannot also exert social influence. Group cohesion or similarity can manifest in many ways, such as appearance, language style, identity, and more. We leave this for future work.
@@ 5.4.3. Optimal Number of Agents for Opinion Shifts
In designing multi-agent systems, however, our findings show that simply increasing the number of agents in a system does not necessarily help opinion change - indeed, increasing from three to five agents had the reverse effect of inducing psychological reactance and reducing this change. While the correlation between group size and social pressure is well-known (Asch, 1955), there is scarce literature on the concrete threshold of group numbers needed to trigger psychological reactance (even in traditional human psychology). More research is needed to investigate how many, and under what conditions, a given number of agents will or will not result in resistance rather than conformance to a suggested opinion.
However, regardless of group size, our study suggests the importance of avoiding designs where users feel ”ganged up” upon by the agents. Even though in our study, only the 5-agent group participants clearly reported these feelings, there were also hints of such feelings in the 3-agent group: P51 felt like an outsider ”for a split second”, and P57 ”kind of felt like the odd one out”. Designers should therefore take care to avoid inducing such feelings; possible ways to do so include having the agents express empathy and practice active listening, or having a small portion of the bots side with instead of against the user (a phenomenon similar to minority influence (Gardikiotis, 2011)). These are promising directions for future work.
@@
Given the heightened vulnerability of younger and less educated users, these groups may be at greater risk of manipulation by coordinated AI agents. Platforms and policymakers need to consider not only the regulation of content but also the vulnerability of audiences. Stronger protections, such as age-sensitive safeguards (Sas et al., 2023), educational transparency notices, and digital literacy interventions, may be needed to ensure that the use of persuasive multi-agent systems does not disproportionately impact these more susceptible populations.
@@ 5.5. Limitations & Future Work
Firstly, given the exploratory and foundational nature of our study, we chose contemporary social issues that we considered to be fairly neutral, yet relevant enough in public discourse that participants would be capable of discussing them with the agents. With our study demonstrating the presence of social influence effects in these issues, future studies could explore potentially more contentious areas like health, ethics, or politics. Given that people tend to hold stronger and possibly more polarized opinions on these issues, future work could explore if multi-agent social influence effects still persist in such cases, or whether new effects might emerge.
Secondly, our study was conducted as a one-time experiment lasting less than an hour, which may not fully capture potential long-term shifts in participants’ attitudes. Opinion change is often a dynamic process influenced by various contextual and temporal factors. Additionally, the relatively small sample sizes in each experimental group may have contributed to greater variability in some measures, as reflected in the larger standard deviations and the spread observed in the boxplots. Future research could address these limitations by incorporating longer-term interventions, larger and more diverse samples, and follow-up assessments to better understand how social influence from multiple AI agents unfolds over time and in broader populations.
Thirdly, since social influence can be shaped by cultural norms, the cultural background of participants is important to consider. Our study included participants from predominantly Western, individualistic cultures. In more collectivist societies, where social norms may have a stronger influence, the results could differ. Future work should examine the impact of multi-agent systems on social influence across diverse cultural contexts.
Finally, we only studied the most fundamental form of social group influence, in which all members of the group hold the same opinion. Future work could investigate different scenarios involving groups with mixed opinion compositions, such as one dissenting agent or varying degrees of agreement among agents, to reflect other realistic and more nuanced group dynamics. Additionally, future studies could also implement direct measurements of psychological reactance and social pressure to validate our proposed explanatory mechanisms. Finally, we only tested three group sizes (1, 3, and 5 agents), following past literature studying varying group sizes (e.g. (Beinema et al., 2021; Jiang et al., 2023; Cacioppo and Petty, 1979)). Nonetheless, future studies could investigate other group sizes (e.g., 2-agent and 4-agent) to better capture the nonlinear effects of group size on social influence.
@@ 6. Conclusion
As multi-agent systems become increasingly integrated into daily life, understanding how these agents interact with humans and influence our perceptions and opinions remains an underexplored yet significant issue. To address this gap, we investigated how interacting with varying numbers of agents affects individuals’ perceptions and opinions through a human-agent social discussion experiment. By analyzing quantitative and qualitative data, we found that multi-agent groups led to more substantial shifts in participants’ opinions towards those of the agents (RQ1). Additionally, participants reported a stronger sense of normative influence when interacting with multiple agents, especially when the agents disagreed with them. We identified potential factors contributing to this social influence, including perceived pressure to conform, motivated by a desire to affiliate with the group of agents (RQ2). Furthermore, demographic analysis indicated that younger participants were more likely to be influenced by multi-agents, potentially due to their higher susceptibility to social pressure and norms compared to older individuals (RQ3). Overall, these findings suggest that agents can function as a social group, exerting social influence on human participants similar to that of human groups. This insight extends the existing CASA framework by demonstrating how interactions with multiple agents can create a sense of group identity. Our study also underscores the potential of multi-agent systems for behavior intervention applications, offering design recommendations while cautioning against the risks and potential harms associated with their misuse.
Acknowledgements.
This research was supported by National University of Singapore (HSS SEED FUND CR 2024) and the Google SEA research award. We thank all reviewers’ comments and suggestions to help polish this paper.
@@ A.1.1. Quantitative
Table 1. Summary of Users’ General Impressions Across Experimental Conditions
Item Scenario 1 Agent 3 Agents 5 Agents H/F p
Mean STD Mean STD Mean STD
Understanding
Topic 1 5.35 1.84 5.45 1.06 5.30 1.37 1.28 0.53
@@
Table 1 presents users’ general impressions of the agent(s), using the same analysis methods described in Section 3.5.
@@ A.1.2. Qualitative
A.2. Control Variables
Table 2 reports the results for the control variables, which showed no significant differences across groups.
Table 2. Summary of Control Variables Across Experimental Conditions
Item Type 1 Agent 3 Agents 5 Agents H/F p
Mean STD Mean STD Mean STD
Topic 1 Expertise
Experience 0.16 0.37 0.18 0.39 0.31 0.47 2.28 0.32
Familiarity 2.65 0.66 2.91 0.77 2.79 0.62 3.66 0.16
Frequency 4.94 1.03 5.39 0.56 5.03 1.09 3.77 0.15
Topic 2 Expertise
Experience 0.74 0.44 0.94 0.24 0.79 0.41 4.68 0.10
Familiarity 3.23 0.80 3.42 0.66 3.34 0.72 0.88 0.64
Frequency 3.81 1.51 4.52 1.15 4.55 1.18 4.88 0.09
Conformity
Positive 5.87 0.73 5.37 1.06 5.25 1.11 4.80 0.10
Negative 3.18 1.11 3.68 1.26 3.61 1.49 1.35 0.27
Compliance
Positive 5.45 1.07 4.86 1.35 5.14 1.13 1.97 0.15
Negative 4.40 1.18 4.37 1.23 4.56 0.94 0.25 0.78
AI Acceptance - 4.18 1.55 4.71 1.32 4.72 1.05 2.12 0.35
A.3. Survey Items
A.3.1. RQ1 - Opinion Change